Prev Next

SW&IT Meeting Minutes 2022-02-28

Attendees: Brandon, David, Glenn, Matthieu, Paul, Stephen, Xianyu, Yang

Outlook

Labor Reporting

Report auto-generated every Monday at 6:30am, Do not edit this table, Update your entries in smartsheets.

20220221
LR20220221 - Bec, Matthieu 40.0
LR20220221 - Mechtley, Brandon 0.0
LR20220221 - Eychaner, Glenn 40.0
LR20220221 - Edmeades, David 24.0
LR20220221 - Grenz, Paul 20.0
LR20220221 - Hooper, Stephen 0.0
LR20220221 - Zhang, Yang 40.0
LR20220221 - Zhang, Xianyu 40.0

News from management

Previous Action Items

Owner Action Due Date Completion Date

Open topics

Team reports

MB

  • Last week
    • OC, Management, ATT PM
    • ao-confcalib autmoated rev control - submit PR and deploy
    • LN loadrecmat default fixes
  • This week
    • set_z0 testing
    • LN tests
    • AO on sky efficiency task definitions (scope, resources, timeline, priority) meeting
    • AO SW requirements in the SOUL commissioning report and in the verification matrix meeting
    • SHARK-NIR: Risk & Hazard Assessment meeting

YZ

  • Previous:
    • Fixed problems in Argos white light source proxy server.
    • LBC bias levels work with Alessandro and Olga (cable swap on Wed and Fri, Bias level changed on Wed, Fri, Sat).
    • Checking and cleaning old MODS images under /lbt/data/new.
    • Working on understanding LBC/DOFPIA behavior and problems with Olga.
    • TMS remote shutter software update.
    • TMS test on Friday.
    • Working with Olga on questions to OSU team related to IT#8567.
    • Drafted archive software request (to be sent on Monday).
  • Next:
    • Remember to send archive software request.
    • Routine TMS software work, and possibly another TMS software test.
    • Continue LBC/DOFPIA study with Olga.
    • Work on cleaning historical data on the MODS data machines.
    • Working with Olga on questions to OSU team related to IT#8567.
    • OSS patch release and proxy setup (maybe a final test)??

PSG

  • Last week
    • GCS Work, fix completed, pull request made
    • NOMIC work, still not complete.
    • SHARK-NIR landing page updated and installed
  • This week
    • Missing header items in NOMIC.
    • Finish NOMIC code changes for header.
    • Look into other GCS issutraks.

GE

  • Last week: IssueTrak week
    • IT 8578 Spent two days digging through logs to find the source of the issue (the AO commit() locking the DataDictionary) and fix. Deployed LUCI 2.14.5.
    • IT 8579 Discovered that the rotator speed was triggering a change from clockwise to counter-clockwise in the PFC calculation. Next action to SciOps.
    • IT 8570 Discovered that turning off flexure correction was turning off data subwindows. Changed so that an explicit empty subwindow (null position values) is required to turn off data subwindows. Deployed LUCI 2.14.6.
    • IT 6818 Fixed bugs in the dynamic amplifier adjustment in the LUCI2 RTD. Monitoring.
    • IT 8584 Investigated but could find no reason the images weren't displayed. Monitoring.
    • IT 8494 Added an explicit change to XY coordinates after receiving more info from SciOps regarding the issue. Monitoring.
    • Also revised or added comments to a number of other ITs.
  • This week
    • Keep testing the new RTD and fix minor bugs as they come up.
    • Restart work on the observer GUI.
    • Clean up the scripts in lcsp/bin on luci.luci and in stow on observer machines. They have drifted significantly from what is checked into GitHub.
    • Investigate further the black-background (miscolored) LUCI windows; D.Thompson found another instance of the problem using ssh from obs4 to luci.luci.
    • Examine the other LUCI GUIs to check for GUI modifications from threads other than the Event Dispatch Thread.

BM

  • Last week:
    • re-ordered gain ramp, tested with LBTI, SOUL
    • new issues:
      • daytime.conf not being read consistently
      • if autogain is interrupted by loopfault, its gains should be reset to 1.
    • ao sw test wiki page
    • confcalib meeting, ln planning meeting
    • argos testing conf file kerfuffle
  • This week:
    • set_z0 test/planning.
    • LBTI offsetting/pause/resume bug.
    • ICE dispatch errors.
    • Debug housekeeper fix

SH

  • VMware PERC cards
  • Reconfigure AMD machine
  • AO switch configuration with LeRoy
  • Moving data from old to new Synology
  • David Carroll monitor
  • SRC, IT meeting

next:

  • Moving data from old to new Synology
    • start LBTI
    • Telemetry RPM
  • VMware installations
    • licenses
  • Upgrade 507 computer
Topic revision: r8 - 28 Feb 2022, BrandonMechtley
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback